OpenStudy (eric_d):

Differentiate following with respect to x using logaritms. 25) (2-x^3)^1/3 / x

OpenStudy (zzr0ck3r):

\(y=\frac{(2-x^3)^\frac{1}{3}}{x}\\\ln(y)=\ln(\frac{(2-x^3)^\frac{1}{3}}{x})=\frac{1}{3}\ln(2-x^3)-\ln(x)\)

OpenStudy (zzr0ck3r):

you with me?

OpenStudy (eric_d):

Srry. OS lag..

OpenStudy (zzr0ck3r):

now we can differentiate term by term, and use explicit differentiation \(\frac{1}{y}\frac{dy}{dx}=\frac{1}{3}\frac{-3x^2}{2-x^3}-\frac{1}{x}\\\frac{dy}{dx}=y(\frac{1}{3}\frac{-3x^2}{2-x^3}-\frac{1}{x})\) now sub in for \(y\) \(\frac{dy}{dx}=(\frac{(2-x^3)^\frac{1}{3}}{x})(\frac{1}{3}\frac{-3x^2}{2-x^3}-\frac{1}{x})\)

OpenStudy (zzr0ck3r):

can you simplify?

OpenStudy (eric_d):

will try

OpenStudy (eric_d):

For the 2nd part.. like this ?